OpenStudy (calculusxy):

Check exponents problem. WILL MEDAL!!! I just need someone who will check some exponents problem. I solved the problems, just need someone to double check.

OpenStudy (calculusxy):

\[x^5 \times x^3 = x^8\]

OpenStudy (calculusxy):

@agent0smith @mathmale

OpenStudy (skullpatrol):

Correct

OpenStudy (calculusxy):

\[\frac{6^5}{6^3} = 6^2 = 36\]

OpenStudy (skullpatrol):

Correct

OpenStudy (calculusxy):

\[\frac{x^4}{x^7} = x^{-3} = \frac{1}{x^3}\]

OpenStudy (anonymous):

yes, multiplying with exponents all you do is add the exponents then dividing with variables would be subtracting

OpenStudy (skullpatrol):

Correct

OpenStudy (calculusxy):

\[(y^4)^3 = y^{12}\]

OpenStudy (skullpatrol):

Correct

OpenStudy (calculusxy):

\[(x^2y)^4 = x^8y^4\]

OpenStudy (skullpatrol):

Correct

OpenStudy (calculusxy):

\[\frac{8x^5}{4x^2} = 2x^3\]

OpenStudy (skullpatrol):

Correct

OpenStudy (calculusxy):

\[\frac{x^5y^6}{xy^2} = x^4y^4\]

OpenStudy (skullpatrol):

Correct

OpenStudy (calculusxy):

\[(\frac{5x^3y}{20xy^3})^4 = \frac{5^4x^{12}y^4}{20^4x^4y^{20}} = \frac{x^8}{4y^{16}}\]

OpenStudy (calculusxy):

\[7^{-2} = \frac{1}{49}\]

OpenStudy (calculusxy):

\[\frac{1}{x^{-5}} = x^5\]

OpenStudy (calculusxy):

\(x^{-6} = \frac{1}{x^6}\)

OpenStudy (skullpatrol):

The last three are correct, but the complicated one is not.

OpenStudy (calculusxy):

How do we do that then?

OpenStudy (skullpatrol):

Look at the exponent of the y in the denominator

OpenStudy (calculusxy):

Oh!

OpenStudy (calculusxy):

It was supposed to be \(y^{12}\) right?

OpenStudy (skullpatrol):

Yes

OpenStudy (calculusxy):

Wait. I just looked at the original problem. It was supposed to be y^5. I am sorry.

OpenStudy (skullpatrol):

:'(

OpenStudy (calculusxy):

I am sorry :'( But would that answer be correct?

OpenStudy (calculusxy):

Meaning my original answer.

OpenStudy (calculusxy):

@skullpatrol

OpenStudy (skullpatrol):

You should have 4^4 in the denominator

OpenStudy (calculusxy):

We would need to carry the exponent of 4?

OpenStudy (skullpatrol):

Yes

OpenStudy (calculusxy):

Okay. So my answer would be \(\large \frac{x^8}{256y^{16}}\)

OpenStudy (skullpatrol):

Correct the question please.

OpenStudy (calculusxy):

What do you mean?

OpenStudy (skullpatrol):

y^5

OpenStudy (calculusxy):

\[\large (\frac{ 5x^3y }{ 20xy^5})^4 = \frac{ x^8 }{ 256y^{16} }\]

OpenStudy (skullpatrol):

Correct :D

OpenStudy (calculusxy):

Okay moving on :)

OpenStudy (calculusxy):

\[x^9 \times x^{-7} = x^2\]

OpenStudy (skullpatrol):

Correct

OpenStudy (calculusxy):

\[\frac{ x^{-1} }{ x^{-8} } = x^7\]

OpenStudy (skullpatrol):

Correct

OpenStudy (calculusxy):

\[\frac{ x^{-4} }{ x^{-9} } = x^5\]

OpenStudy (skullpatrol):

Correct

OpenStudy (calculusxy):

\[(2x^3y^{-3})^{-2} = \frac{y^6}{4x^6}\]

OpenStudy (skullpatrol):

Correct

OpenStudy (calculusxy):

\[(4x^4y^{-4})^3 = \frac{164x^{12}}{y^{12}}\]

OpenStudy (calculusxy):

sorry i meant 64

OpenStudy (calculusxy):

\[\frac{ 64x^{12} }{ y^{12} }\]

OpenStudy (skullpatrol):

Correct

OpenStudy (calculusxy):

\[(5x^2y)(2x^4y^{-3}) = \frac{10x^6}{y^3}\]

OpenStudy (skullpatrol):

y^3 in the denominator?

OpenStudy (calculusxy):

Yes. Why?

OpenStudy (skullpatrol):

What about the y in the numerator

OpenStudy (calculusxy):

What do you mean?

OpenStudy (skullpatrol):

5x^2y

OpenStudy (calculusxy):

It's supposed to be a multiplication problem. I don't see why \(y\) needs to be in the numerator.

OpenStudy (skullpatrol):

$$\Huge 5x^2y$$

OpenStudy (calculusxy):

I still don't understand. Please make it clear.

OpenStudy (calculusxy):

I just checked it out with Wolfram Alpha. They say that the denominator needs to have \(y^2\)?

OpenStudy (skullpatrol):

$$\Huge y^{-3}=\dfrac{1}{y^3}$$

OpenStudy (calculusxy):

Oh! \(\large y^1 \times y^{-3} = y^{1 + (-3)} = y^{-2} = \frac{1}{y^2}\)

OpenStudy (skullpatrol):

Yes

OpenStudy (calculusxy):

OK got it! Thank you so much for your time and help! :)

OpenStudy (skullpatrol):

Thanks for trying to learn :-)